OpenStudy (anonymous):

A stable population P of mosquitoes (measured in thousands), is observed to suddenly undergo a period of acceleration, given by a(t) = 2 − [1/(t + 4)] after t days. If the population was initially 1000 mosquitoes, determine the function P(t), and then find the population after 21 days. I got: v(t)= 2t - (2)(t+4)^(1/2) + C P(t) = t^2 - (4/3)(t+4)^(3/2) +Ct + K K=1000 But I’m not sure how to get C

OpenStudy (anonymous):

\[\int\limits_{}^{}a(t)=2-[\frac{ 1 }{ t+4 }]\]\[v(t)=2t-\ln(t+4) +C\] \[P(t)=t ^{2}-[(t+4)\ln(t+4)-(t+4)] +ct+k\] I think this is how the integration should go

OpenStudy (anonymous):

it's 1/sqrt(t+4)

OpenStudy (anonymous):

And I still need to find c

OpenStudy (anonymous):

sorry